The equation for photon energy, E, is

E=hcλ

where h = 6.626×10−34 J⋅s (Planck's constant) and c = 2.99×108 m/s (the speed of light).

What is the wavelength, λ, of a photon that has an energy of E = 3.39×10−19 J ?

Express your answer numerically in meters.
Answer #1

We have,

Energy of the photon, E = hc/

Where, h is the Planck’s constant and c is thewavelength of light

= hc/E = (6.626×10-34 J⋅s) x (2.99×108 m/s)/(3.39 x 10-19 J) = 5.84 x 10-7 m

Wavelength of the photon = 5.84 x 10-7 m